# Taming High-Dimensional Dynamics: Learning Optimal Projections onto Spectral Submanifolds

### v1

@@ Abstract
High-dimensional nonlinear systems pose considerable challenges for modeling and control across many domains, from fluid mechanics to advanced robotics. Such systems are typically approximated with reduced order models, which often rely on orthogonal projections, a simplification that may lead to large prediction errors. In this work, we derive optimality of fiber-aligned projections onto spectral submanifolds, preserving the nonlinear geometric structure and minimizing long-term prediction error. We propose a computationally tractable procedure to approximate these projections from data, and show how the effect of control can be incorporated. For a 180-dimensional robotic system, we demonstrate that our reduced-order models outperform previous state-of-the-art approaches by up to fivefold in trajectory tracking accuracy under model predictive control.
@@ I Introduction
We address the challenge of modeling and controlling high-dimensional nonlinear systems, prevalent in fields like fluid mechanics and robotics, where computational complexity necessitates reduced-order models (ROMs). Traditional ROMs, typically based on orthogonal projections like Proper Orthogonal Decomposition (POD) [berkooz2003proper], often fail to accurately capture nonlinear dynamics critical for reliable long-term prediction and control [cenedese2022data-driven].
Reducing the dynamics to Spectral Submanifolds (SSMs) offers a rigorous alternative for capturing dominant slow dynamics in nonlinear systems [haller2016nonlinear]. In general, the system’s trajectory will not lie exactly on an SSM, requiring a mapping from off-manifold states onto on-manifold states. Prior data-driven SSM methods [alora2025discovering, AloraCenedeseEtAl2023, kaundinya2025data-driven, alora2023robust] carry out this mapping via orthogonal projections onto the tangent space of the SSM. While often sufficient, such orthogonal projections neglect the system’s stable fiber geometry, leading to inaccuracies in capturing transient behaviors and thus limiting ROM performance in predictive control applications.
@@
Outline: Section II reviews related work. Section III formally defines slow SSMs and their stable foliations. Section IV establishes the optimality of fiber-aligned projections and introduces a data-driven approach to learning the projection operator. Section V applies SSMs to control and provides numerical validations and comparisons, with conclusions in Section VI. For brevity, we refer the reader to proofs in Sections -C, -B and -A.
@@ II Related work
Linear projection-based methods, such as POD, often utilize oblique projections (i.e., Petrov-Galerkin) to address the challenges of reducing systems which exhibit significant nonnormality, and transient growth (see [benner_reviewPOD_2015] for a comprehensive review). More rigorous alternatives like SSM-based model reduction also utilize oblique projections, but unlike the data-driven projections in POD, these are inferred directly from a modal analysis of the slow-fast splitting of the governing equations [jain_2022_SSMTool].
While data-driven SSMs have shown state-of-the-art performance in various applications spanning fluids and soft robot control [cenedese2022data-driven, alora2025discovering, kaundinya2025data-driven] the fast and slow subspaces of the systems in these applications are often non-orthogonal at the fixed point from which the SSM emanates [bettini_oblique_2025]. To tackle this for 2D SSMs, Bettini et al. [bettini_oblique_2025] optimize an oblique projection using backbone data from mechanical vibration experiments. This local approximation of the fast stable invariant foliation is a more efficient alternative to learning global, nonlinear fiber approximations [szalai2023_ISF], since this demands on significant amounts of data.
Alternatively, autoencoder-based reduction strategies project dynamics onto slow manifolds, determining their dimension by minimizing training error [lee_2020_autoencoder, Champion_2019_autoencoder, fresca_2021_autoencoder]. These methods, however, often project orthogonally onto the target structure, introducing errors. Otto et al. [otto2023learning] refine this by incorporating nonlinear oblique projections, learning an optimal encoder to minimize the gap between projected and reduced dynamics. While effective, this approach requires knowledge of the governing equations and is data-intensive, also rendering it impractical for hardware implementations.
Inspired by the effectiveness of optimal linear oblique projections in experimental settings [bettini_oblique_2025] and the broad generality of nonlinear projections [otto2023learning], we propose a method that learns an optimal linear oblique projection onto SSMs directly from data. This approach strikes a valuable trade-off, blending the practicality of [bettini_oblique_2025] with the versatility of [otto2023learning], making it well-suited for diverse applications, including high-dimensional systems like soft robots. We further extend this method to nonlinear systems using more recent results on Normally Attracting Invariant Manifolds (NAIMs) [eldering2013normally, eldering2018global].
@@ III Structure of a slow SSM
Assumption 1 is a standard smoothness assumption [Lorenz2005, Cannarsa2006] that guarantees the existence and uniqueness of solutions to the ODEs in (1). By multiplying $\mathbf{f}$ with a smooth cutoff function whose arbitrarily large support contains states of interest, the Lipschitz continuity assumptions are always satisfied if $\mathbf{f}\in C^{2}$.
@@ III-A Slow SSM
This definition is a restatement of the slow SSM definition in [haller2016nonlinear], combined with the fact that a slow stable SSM with a single fixed point is a uniform NAIM [eldering2013normally, hirsch1970invariant]. The low-dimensionality of $\mathcal{M}$ manifests due to the uniform attraction condition (3), which ensures that the contraction and expansion rates normal to $\mathcal{M}$ dominate those within it. As a result, the long-term dynamics of Equation 1 are effectively confined to $\mathcal{M}$, making it a natural candidate for reduced-order modeling.
@@ III-B Local Coordinates and Flow Near $\mathcal{M}$
To study the dynamics near a compact slow spectral submanifold $\mathcal{M}$, we construct local coordinate systems using standard manifold theory [wiggins1994normally, fenichel1974asymptotic]. Because our analysis is local and focused on behavior close to $\mathcal{M}$, we now introduce a suitable neighborhood around it. Specifically, we define the tubular neighborhood of $\mathcal{M}$, denoted by $\mathcal{T}^{\varepsilon}$, by first considering the normal $\varepsilon$-neighborhood
@@
$\varphi_{t}(\mathbf{x}_{r},\mathbf{x}_{n})\vcentcolon=\bm{\sigma}\circ\mathbf{h}^{-1}\circ\mathbf{F}^{t}\circ\mathbf{w}(\mathbf{x}_{r},\mathbf{x}_{n}).$ (7)
Hence, $\varphi_{t}(\mathbf{x}_{r},\mathbf{x}_{n})$ represents how the reduced coordinates evolve under the flow.
@@ III-C Foliations of a Slow SSM
We say that $\mathbf{x}$ is on the stable fiber of the base point $\mathbf{p}$ when $\mathbf{x}\in\mathcal{F}(\mathbf{p})$. The stable foliation, $\mathcal{W}_{\mathrm{loc}}^{s}(\mathcal{M})$ is then given by
@@
We now state the invariant foliation theorem [wiggins1994normally, fenichel1974asymptotic] which outlines the properties of the stable structure around $\mathcal{M}$.
@@
Let $\mathcal{M}$ be a compact $C^{r}$-smooth slow SSM of system (1). Then there exists a ${n_{f}}$-parameter family of $({n_{f}}-n)$-dimensional stable fibers $\mathcal{W}_{\mathrm{loc}}^{s}(\mathcal{M})$ satisfying:
@@
By definition, a slow SSM is a NAIM [eldering2013normally]. We then invoke results from Theorem 1 in [eldering2018global] that guarantee the existence of a stable foliation. Since a NAIM is a special case of a normally hyperbolic invariant manifold, the properties of the stable foliation as stated in [wiggins1994normally] apply. ∎
@@ IV Optimal projection operator
In this section, we show that the fibers implicitly characterize the optimal projection operator onto the slow SSM. We first prove that projecting initial conditions onto their corresponding fiber’s base point minimizes the integrated error. We then show the form of the fiber map, $f^{s}(\mathbf{p})$, up to first order. This allows us to derive a linear oblique projection onto $\mathcal{M}$, for which we propose an efficient data-driven learning algorithm.
@@ IV-A Optimality of the Base Point
[…]
@@
[…]
@@
[…]
@@
[…]
@@
where the first inequality is due to Lemma 1 and the second is due to Lemma 2. Plugging this into (17) and integrating over $[0,T_{1}]$ gives
@@
Hence, choosing $\delta$ small enough so that
@@
ensures the desired inequality, concluding the proof. ∎
@@ IV-B Form of Projection Operator
$\bm{\nu}(\mathbf{x})$ :$\vcentcolon=\bm{\sigma}\circ\mathbf{h}^{-1}(\mathbf{x}),$ (21)
$\bm{\omega}(\mathbf{x}_{r})$ :$\vcentcolon=\mathbf{w}(\mathbf{x}_{r},\mathbf{0}),$
@@
where $\mathbf{V}_{\text{E}}=\begin{bmatrix}{\mathbf{I}}_{n\times n}^{\top}&\mathbf{0}_{n\times{n_{f}}-n}^{\top}\end{bmatrix}^{\top}$ and $\mathbf{V}_{\text{opt}}=\begin{bmatrix}{\mathbf{I}}_{n\times n}&-\mathbf{V}_{\text{opt}}\end{bmatrix}$.
@@
By Theorem 1, $f^{s}(\cdot;\mathbf{p})$ is a $C^{r}$-smooth surface with the key property that its dependence on the base point $\mathbf{p}$ is also $C^{r}$ smooth. Expanding $f^{s}$ in $\mathbf{x}_{n}$ around $\bar{\mathbf{x}}_{n}=\mathbf{0}$ for fixed $\mathbf{p}$, we have
$\mathbf{x}_{r}$ $=f^{s}(\mathbf{x}_{n};\mathbf{p})=f^{s}(\mathbf{0};\mathbf{p})+\mathbf{D}_{\mathbf{x}_{n}}f^{s}(\mathbf{0};\mathbf{p})\mathbf{x}_{n}+\mathcal{O}(\left\|\mathbf{x}_{n}\right\|^{2})$ (25)
$=\bar{\mathbf{x}}_{r}+\mathbf{D}_{\mathbf{x}_{n}}f^{s}(\mathbf{0};\mathbf{p})\mathbf{x}_{n}+\mathcal{O}(\left\|\mathbf{x}_{n}\right\|^{2})$
$=\bar{\mathbf{x}}_{r}+\mathbf{D}_{\mathbf{x}_{n}}f^{s}\big{(}\mathbf{0};(\bar{\mathbf{x}}_{r},\mathbf{g}(\bar{\mathbf{x}}_{r}))\big{)}\mathbf{x}_{n}+\mathcal{O}(\left\|\mathbf{x}_{n}\right\|^{2}),$
where the first equality is due to the definition of a fiber (8) and the second equality is due to the SSM being constructed as a graph, where $\mathbf{p}=(\bar{\mathbf{x}}_{r},\mathbf{g}(\bar{\mathbf{x}}_{r}))$.
Expanding $f^{s}(\cdot;(\bar{\mathbf{x}}_{r},\mathbf{g}(\bar{\mathbf{x}}_{r})))$ around $\bar{\mathbf{x}}_{r}=\mathbf{0}$ we have
$f^{s}\big{(}\mathbf{x}_{n};(\bar{\mathbf{x}}_{r},\mathbf{g}(\bar{\mathbf{x}}_{r}))\big{)}=$ $~{}f^{s}\big{(}\mathbf{x}_{n};(\mathbf{0},\mathbf{0})\big{)}$ (26)
$+\mathbf{D}_{\mathbf{x}_{r}}f^{s}(\mathbf{x}_{n};(\mathbf{0},\mathbf{0}))\mathbf{x}_{r}$
$+\mathcal{O}(\left\|\bar{\mathbf{x}}_{r}\right\|^{2}),$
where we invoke tangency of the manifold at the origin, hence $\mathbf{D}_{\mathbf{x}_{r}}\mathbf{g}(\mathbf{0})=\mathbf{0}$ and $\mathbf{g}(\mathbf{0})=\mathbf{0}$. Combining (25) and (26), yields the following
$\begin{bmatrix}\bar{\mathbf{x}}_{r}\\ \mathbf{0}\end{bmatrix}=\begin{aligned} \begin{bmatrix}\mathbf{x}_{r}-\mathbf{V}_{0}\mathbf{x}_{n}-\mathbf{V}_{1}(\bar{\mathbf{x}}_{r})\mathbf{x}_{n}\\ \mathbf{0}\end{bmatrix},\end{aligned}$
where $\mathbf{V}_{0}=\mathbf{D}_{\mathbf{x}_{n}}f^{s}(\mathbf{0};\mathbf{0})$ and $\mathbf{V}_{1}(\bar{\mathbf{x}}_{r})=\mathbf{D}_{\mathbf{x}_{n}}\mathbf{D}_{\mathbf{x}_{r}}f^{s}(\mathbf{0};\mathbf{0})$ is a constant third order tensor in $\mathbb{R}^{n\times({n_{f}}-n)\times n}$. We take a zeroth-order approximation and define the candidate projection operator
$\Pi(\mathbf{x})=\begin{bmatrix}\bar{\mathbf{x}}_{r}\\ \mathbf{0}\end{bmatrix}=\underbrace{\begin{bmatrix}{\mathbf{I}}_{n\times n}\\ \mathbf{0}_{{n_{f}}-n\times n}\end{bmatrix}}_{\mathbf{V}_{\text{E}}}\underbrace{\begin{bmatrix}{\mathbf{I}}_{n\times n}&-\mathbf{V}_{0}\end{bmatrix}}_{\mathbf{V}_{\text{opt}}}\begin{bmatrix}\mathbf{x}_{r}\\ \mathbf{x}_{n}\end{bmatrix}\end{aligned}.$
Note that $\mathbf{V}_{\text{E}}$ and $\mathbf{V}_{\text{opt}}$ are not unique since we can always apply a similarity transform to these matrices. ∎
This proposition shows that nearby fibers are approximately pointing in the same direction as the stable fiber at the equilibrium point. To refine the current approximation to a higher order, we replace the constant $\mathbf{V}_{0}$ with $\mathbf{V}_{0}+\mathbf{V}_{1}\mathbf{x}_{r}$, where $\mathbf{x}_{r}$ represents the reduced coordinate of the fiber base point. Although this introduces additional unknowns through $\mathbf{V}_{1}$, it allows for a more accurate representation of the fiber’s behavior near the equilibrium. In this work, we focus on a zeroth-order approximation of the fiber attached to the stable fixed-point at the origin, i.e.,
$\mathbf{x}_{r}=\bar{\mathbf{x}}_{r}+\mathbf{V}_{0}\mathbf{x}_{n},$ (27)
where, with slight abuse of notation, we denote the point $\mathbf{x}=(\mathbf{x}_{r},\mathbf{x}_{n})$ as a point in the neighborhood of the manifold. [^4] Using local coordinates on the manifold, where $(\mathbf{x}_{r},\mathbf{x}_{n})=(\bm{\sigma}\times\bm{\tau})(\mathbf{x})$.
@@
where $\mathbf{A}_{TT}\in\mathbb{R}^{n\times n}$ acts on the tangent directions, $\mathbf{A}_{TN}\in\mathbb{R}^{n\times{n_{f}}-n}$ accounts for first-order coupling from normal to tangent direction, and $\mathbf{A}_{NN}\in\mathbb{R}^{{n_{f}}-n\times{n_{f}}-n}$ acts on the normal direction. To show that $(\mathbf{V}_{0}\mathbf{x}_{n},\mathbf{x}_{n})\in\mathbf{E}^{s}_{\mathbf{0}}$, we need to show that the subspace
@@
which implies we must find $\mathbf{V}_{0}$ which solves the Sylvester equation below
@@
There exists a unique solution to (30) iff $\text{spect}(\mathbf{A}_{TT})\cap\text{spect}(\mathbf{A}_{NN})=\emptyset$ [golub2013matrix], which is guaranteed in our setting due to the spectral gap between $E$ and $\mathbf{E}^{s}_{\mathbf{0}}$. ∎
The projection operator $\Pi$ is valid only near the fixed point, where its linear approximation captures transient behavior effectively. In the following section we will learn this projection operator across the phase space of the system. It is important to note that learning $\Pi$ from data far from this point minimizes average error by fitting prevalent transients, but not long term error as in (15). While learning true nonlinear fibers will minimize (15), it is typically data intensive to do so. Though not exact, our method balances accuracy and practicality, providing a feasible alternative to the ideal.

### v2

@@ Abstract
High-dimensional nonlinear systems pose considerable challenges for modeling and control across many domains, from fluid mechanics to advanced robotics. Such systems are typically approximated with reduced-order models, which often rely on orthogonal projections, a simplification that may lead to large prediction errors. In this work, we derive optimality of fiber-aligned projections onto spectral submanifolds, preserving the nonlinear geometric structure and minimizing long-term prediction error. We propose a data-driven procedure to learn these projections from trajectories and demonstrate its effectiveness through a 180-dimensional robotic system. Our reduced-order models achieve up to fivefold improvement in trajectory tracking accuracy under model predictive control compared to the state of the art.
@@ I Introduction
We address the challenge of modeling and controlling high-dimensional nonlinear systems, prevalent in fields like fluid mechanics and robotics, where computational complexity necessitates reduced-order models (ROMs). Traditional ROMs, typically based on orthogonal projections like Proper Orthogonal Decomposition (POD) [1], often fail to accurately capture nonlinear dynamics critical for reliable long-term prediction and control [2].
Reducing the dynamics to Spectral Submanifolds (SSMs) offers a rigorous alternative for capturing dominant slow dynamics in nonlinear systems [3]. In general, the system’s trajectory will not lie exactly on an SSM, requiring a mapping from off-manifold states onto on-manifold states. Prior data-driven SSM methods [4, 5, 6, 7] carry out this mapping via orthogonal projections onto the tangent space of the SSM. While often sufficient, such orthogonal projections neglect the system’s stable fiber geometry, leading to inaccuracies in capturing transient behaviors and thus limiting ROM performance in predictive control applications.
@@
Outline: Section II reviews related work. Section III formally defines slow SSMs and their stable foliations. Section IV establishes the optimality of fiber-aligned projections and introduces a data-driven approach to learning the projection operator. Section V applies SSMs to control and provides numerical validations and comparisons, with conclusions in Section VI. For brevity, we refer the reader to proofs in Sections -C, -B, -A and -D.
@@ II Related Work
Linear projection-based methods, such as POD, often utilize oblique projections (i.e., Petrov-Galerkin) to address the challenges of reducing systems which exhibit significant nonnormality, and transient growth (see [8] for a comprehensive review). More rigorous alternatives like SSM-based model reduction also utilize oblique projections, but unlike the data-driven projections in POD, these are inferred directly from a modal analysis of the slow-fast splitting of the governing equations [9].
While data-driven SSMs have shown state-of-the-art performance in various applications spanning fluids and soft robot control [2, 4, 6] the fast and slow subspaces of the systems in these applications are often non-orthogonal at the fixed point from which the SSM emanates [10]. To tackle this for 2D SSMs, Bettini et al. [10] optimize an oblique projection using backbone data from mechanical vibration experiments. This local approximation of the fast stable invariant foliation is a more efficient alternative to learning global, nonlinear fiber approximations [11], since the latter demands significant amounts of data.
Alternatively, autoencoder-based reduction strategies project dynamics onto slow manifolds, determining their dimension by minimizing training error [12, 13, 14]. These methods, however, often project orthogonally onto the target structure, introducing errors. Otto et al. [15] refine this by incorporating nonlinear oblique projections, learning an optimal encoder to minimize the gap between projected and reduced dynamics. While effective, this approach requires knowledge of the governing equations and is data-intensive, also rendering it impractical for hardware implementations.
Inspired by the effectiveness of optimal linear oblique projections in experimental settings [10] and the broad generality of nonlinear projections [15], we propose a method that learns an optimal linear oblique projection onto SSMs directly from data. This approach strikes a valuable trade-off, blending the practicality of [10] with the versatility of [15], making it well-suited for diverse applications, including high-dimensional systems like soft robots. We further extend this method to nonlinear systems using more recent results on Normally Attracting Invariant Manifolds (NAIMs) [16, 17].
@@ III Structure of a Slow SSM
Assumption 1 is a standard smoothness assumption [18, 19] that guarantees the existence and uniqueness of solutions to the ODEs in (1). By multiplying $\mathbf{f}$ with a smooth cutoff function whose arbitrarily large support contains states of interest, the Lipschitz continuity assumptions are always satisfied if $\mathbf{f}\in C^{2}$.
Assumption 2 is well-motivated as most engineering systems are stable at some equilibrium configuration. Although these systems naturally return to equilibrium when unforced, we require control to make them follow desired trajectories away from this stable point.
@@ III-A Slow SSM
This definition is a restatement of the slow SSM definition in [3], combined with the fact that a slow stable SSM with a single fixed point is a uniform NAIM [16, 20]. The low-dimensionality of $\mathcal{M}$ manifests due to the uniform attraction condition (3), which ensures that the contraction and expansion rates normal to $\mathcal{M}$ dominate those within it. As a result, the long-term dynamics of Equation 1 are effectively confined to $\mathcal{M}$, making it a natural candidate for reduced-order modeling.
@@ III-B Local Coordinates and Flow Near $\mathcal{M}$
To study the dynamics near a compact slow spectral submanifold $\mathcal{M}$, we construct local coordinate systems using standard manifold theory [21, 22]. Because our analysis is local and focused on behavior close to $\mathcal{M}$, we now introduce a suitable neighborhood around it. Specifically, we define the tubular neighborhood of $\mathcal{M}$, denoted by $\mathcal{T}^{\varepsilon}$, by first considering the normal $\varepsilon$-neighborhood
@@
$\varphi_{t}(\mathbf{x}_{r},\mathbf{x}_{n})=\bm{\sigma}\circ\mathbf{h}^{-1}\circ\mathbf{F}^{t}\circ\mathbf{w}(\mathbf{x}_{r},\mathbf{x}_{n}),$ (7)
where $\mathbf{h}^{-1}$ returns a tuple $(\mathbf{p},\mathbf{v})$ and $\bm{\sigma}$ is applied only to $\mathbf{p}$. Hence, $\varphi_{t}(\mathbf{x}_{r},\mathbf{x}_{n})$ represents how the reduced coordinates evolve under the flow.
@@ III-C Foliations of a Slow SSM
We say that $\mathbf{x}$ is on the stable fiber of the base point $\mathbf{p}$ when $\mathbf{x}\in\mathcal{F}(\mathbf{p})$. The stable foliation, $\mathcal{W}_{\mathrm{loc}}^{s}(\mathcal{M})$, is then given by
@@
We now state the invariant foliation theorem [21, 22] which outlines the properties of the stable structure around $\mathcal{M}$.
@@
Let $\mathcal{M}$ be a compact $C^{r}$-smooth slow SSM of system (1), such that $r\in\mathbb{N}^{+}\cup\{\infty,a\}$ where $a$ refers to analytic. Then there exists an $n_{f}$}-parameter family of $({n_{f}}-n)$-dimensional stable fibers $\mathcal{W}_{\mathrm{loc}}^{s}(\mathcal{M})$ satisfying:
@@
By definition, a slow SSM is a NAIM [16]. We then invoke results from Theorem 1 in [17] that guarantee the existence of a stable foliation. Since a NAIM is a special case of a normally hyperbolic invariant manifold, the properties of the stable foliation as stated in [21] apply. ∎
@@ IV Optimal Projection Operator
In this section, we show that the fibers implicitly characterize the optimal projection operator onto the slow SSM. We first prove that projecting initial conditions onto their corresponding fiber’s base point minimizes the integrated error. A zeroth-order approximation of $f^{s}(\mathbf{p})$ allows us to derive a linear oblique projection onto $\mathcal{M}$, for which we propose a data-driven learning algorithm.
@@ IV-A Optimality of the Base Point
$\mathbf{p}=\operatorname*{arg\,min}_{\mathbf{p}^{\prime}}\int_{0}^{\infty}\|\mathbf{F}^{t}(\mathbf{x})-\mathbf{F}^{t}(\mathbf{p}^{\prime})\|\,dt.$ (15)
@@
$\int_{0}^{\infty}\left\|\mathbf{F}^{t}(\mathbf{x})-\mathbf{F}^{t}(\mathbf{p})\right\|dt<\int_{0}^{\infty}\left\|\mathbf{F}^{t}(\mathbf{x})-\mathbf{F}^{t}(\mathbf{p}^{\prime})\right\|dt.$ (16)
@@
$M=\inf_{\mathbf{p}^{\prime}\in\{\hat{\mathbf{p}}:\left\|\mathbf{p}-\hat{\mathbf{p}}\right\|<\gamma_{3}\}}\int_{T_{1}}^{\infty}\left\|\mathbf{F}^{t}(\mathbf{x})-\mathbf{F}^{t}(\mathbf{p}^{\prime})\right\|dt>0.$
@@
$\int_{0}^{T_{1}}\hskip-4.62497pt\Big{(}\left\|\mathbf{F}^{t}(\mathbf{x})-\mathbf{F}^{t}(\mathbf{p})\right\|-\left\|\mathbf{F}^{t}(\mathbf{x})-\mathbf{F}^{t}(\mathbf{p}^{\prime})\right\|\Big{)}dt$ (17)
$\leq\int_{0}^{T_{1}}\hskip-4.62497pt\Big{(}\left\|\mathbf{F}^{t}(\mathbf{p})-\mathbf{F}^{t}(\mathbf{p}^{\prime})\right\|\Big{)}dt\leq(1-\epsilon)M.$
@@
where the first inequality is due to Lemma 1 and the second is due to Lemma 2. Plugging this into (17) and integrating over $0,T_{1}]$ gives
@@
Since $\left\|\bar{\mathbf{x}}_{r}-\mathbf{x}_{r}^{\prime}\right\|<\delta$, we can choose $\delta$ small enough so that
@@
ensures the desired inequality. ∎
@@ IV-B Form of Projection Operator
$\bm{\nu}(\mathbf{x})$ $=\bm{\sigma}\circ\mathbf{h}^{-1}(\mathbf{x}),$ (21)
$\bm{\omega}(\mathbf{x}_{r})$ $=\mathbf{w}(\mathbf{x}_{r},\mathbf{0}),$
@@
where $\mathbf{V}_{\text{E}}=\begin{bmatrix}{\mathbf{I}}_{n\times n}&\mathbf{0}_{n\times({n_{f}}-n)}\end{bmatrix}^{\top}$ and $\mathbf{V}_{\text{opt}}=\begin{bmatrix}{\mathbf{I}}_{n\times n}&-\mathbf{V}_{0}\end{bmatrix}^{\top}$.
@@
The proof Taylor expands the fiber map and is detailed in Appendix -D. ∎
Intuitively, Proposition 1 shows that nearby fibers point approximately in the same direction as the stable fiber at the equilibrium point. Since, in practice, we work in ambient coordinates, $\mathbf{V}_{\text{E}}$ and $\mathbf{V}_{\text{opt}}$ may take a different form and must be identified. The procedure for inferring these matrices from data is detailed in Section IV-C.
@@
where $\mathbf{A}_{TT}\in\mathbb{R}^{n\times n}$ acts on the tangent directions, $\mathbf{A}_{TN}\in\mathbb{R}^{n\times({n_{f}}-n)}$ accounts for first-order coupling from normal to tangent direction, and $\mathbf{A}_{NN}\in\mathbb{R}^{({n_{f}}-n)\times({n_{f}}-n)}$ acts on the normal direction. To show that $(\mathbf{V}_{0}\mathbf{x}_{n},\mathbf{x}_{n})\in\mathbf{E}^{s}_{\mathbf{0}}$, we need to show that the subspace
@@
which implies we must find $\mathbf{V}_{0}$ which solves the Sylvester equation
@@
There exists a unique solution to (27) iff $\text{spect}(\mathbf{A}_{TT})\cap\text{spect}(\mathbf{A}_{NN})=\emptyset$ [23], which is guaranteed in our setting due to the spectral gap between $E$ and $\mathbf{E}^{s}_{\mathbf{0}}$. ∎
The projection operator $\Pi$ is valid only near the fixed point, where its linear approximation captures transient behavior. In the following section we will learn this projection operator across the phase space of the system.